a compound is composed of only C H and O. the combustion of a 0.519g sample of the compound yields 1.24g of CO2 and 0.255g of H2O. What is the empirical formula of this compound
1
Expert's answer
2018-11-05T09:47:09-0500


Answer:

mC /12: mH /1: mO / 16

mC= ?

mH=?

mO=?

mC=12 x 1.24 / 44 = 0.3381 g (C)

mH=2 x 0.255 / 18 = 0.0283 (H)

mO= 0.519 – 0.3381 – 0.0283 = 0.152 g (O)

mC /12: mH /1: mO / 16 = 0.3381 / 12 : 0.0283 / 1 : 0.152 / 16 = 0,028175 : 0.0283 : 0,0095

0,028175 / 0,0095 : 0.0283 / 0,0095: 0,0095 / 0,0095 = 3 :3 : 1

C=3; H=3; O=1; therefore C3H3O x 2 and

Find the final formula C6H6O2 it is Resorcinol